### v1

# Localization behavior in a Hermitian and non-Hermitian Raman lattice

Abstract
We introduce a flexible Raman lattice system for alakaline-earth like atoms to investigate localization behaviors in a quasi-periodic lattice with controllable non-Hermiticity. We demonstrate that critical phases and mobility edges can arise by adjusting spin-dependence of the incommensurate potentials in the Hermitian regime. With non-Hermiticity introduced by spin-selective atom loss, we find the localization behaviour in this system can be suppressed by dissipation. Our work provides insights into interplay between quasi-periodicity and non-Hermitian physics, offering a new perspective on localization phenomena.
I I. Introduction
In the presence of a disorder potential, quantum systems can exhibit exponential localization of wave functions induced by wave interference, a phenomenon known as Anderson localization Anderson (1958). Ultracold atom systems, with their exceptional purity and high controllability, enable detailed exploration of localization phenomena. Recent experiments have investigated Anderson localization in optical lattices with either random disorder Billy et al. (2008) or incommensurate quasi-periodic structures Roati et al. (2008); Lüschen et al. (2018); Xiao et al. (2021).
For ultracold atoms, quasi-periodic optical lattices are created by superimposing two optical lattices with incommensurate wavelengths. In such setups, a notable transition from extended to localized states, commonly referred to as the Anderson transition, can manifest even within one-dimensional configurations Grempel et al. (1982); Aulbach et al. (2004), contrary to uncorrelated disordered systems where such transitions are expected to occur only in dimensions higher than two according to the scaling theory Thouless (1974); Abrahams et al. (1979). These systems also allow for the engineering of mobility edges-boundaries that separate coexisting localized and extended states at discrete energy levels Sarma et al. (1988); Evers and Mirlin (2008); Biddle and Das Sarma (2010); Lüschen et al. (2018); An et al. (2018, 2021). Recent theoretical studies suggest even more complex behavior, including intermediate critical phases and hybrid phases where localized, extended, and critical states coexist in specific one-dimensional quasi-periodic models Wang et al. (2020, 2022).
Recent years have seen growing interest in non-Hermitian physics El-Ganainy et al. (2018); Ashida et al. (2021) across multiple research domains, including electrical circuits Helbig et al. (2020); Ezawa (2019), optics and photonics Lee et al. (2009); Xiao et al. (2020); Zhou et al. (2018); Weidemann et al. (2020), optomechanics Patil et al. (2022), NV centers Zhang et al. (2021); Yu et al. (2022), trapped ions Cao et al. (2023); Zhang et al. (2025), superconducting qubits Chen et al. (2021) and ultracold atom systems Li et al. (2019); Liang et al. (2022); Ren et al. (2022). Non-Hermitian systems display unique features such as exceptional points or rings Zhen et al. (2015); Ren et al. (2022); Liu et al. (2025), novel topological invariants, and the non-Hermitian skin effect Liang et al. (2022); Zhao et al. (2025). The Hatano-Nelson model, a fundamental non-Hermitian topological model, combines asymmetric hopping amplitudes with an on-site random potential Hatano and Nelson (1996). However, a comprehensive understanding of the interplay between disorder and non-Hermiticity, particularly the influences of dissipation on localization behavior, remains largely unexplored Zeng et al. (2017); Zhang et al. (2020); Harter et al. (2016). Furthermore, experimental methods for studying this relationship between disorder and non-Hermiticity remain limited.
In this paper, we introduce a flexible Raman lattice scheme Liu et al. (2013) tailored for 173Yb atoms Song et al. (2018, 2019) or alkaline-earth metal atoms Liang et al. (2023) to investigate localization behaviors. The Raman lattice configuration induces quasi-periodicity using two optical lattices with incommensurate wavelengths, with the weak lattice potential akin to an incommensurate Zeeman potential. By adjusting the frequency of the incommensurate lattice close to the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition, the spin-dependent nature of the lattice potential may lead to the emergence of critical phases or mobility edges within the system. Finally, we generalize the Raman lattice with incommensurate potentials into the non-Hermitian regime with dissipation induced by atom loss Zhao et al. (2025). The analysis reveals that dissipation exerts a significant influence on the localization behaviors within the system, offering a platform to investigate the intricate interplay between disorder potential and non-Hermitian physics.
The subsequent sections of this paper are structured as follows: Sec.II presents the model Hamiltonian of the flexible Raman lattice and outlines its experimental realization scheme. In Sec.III-V, we delve into the impact of the spin-dependency of the incommensurate lattice and dissipation within the system. Our focus will encompass three distinct scenarios: a completely spin-dependent incommensurate lattice, a partially spin-dependent incommensurate lattice, and an incommensurate lattice coupled with dissipation. Finally, Sec.VI offers a concise summary of the findings and conclusions derived from this investigation.
II II. Model Hamiltonian
The Hamiltonian considered in this paper can be written as:
$H=H_{0}+\sum_{i,\sigma=\uparrow,\downarrow}(\delta_{\sigma}^{i}+i\gamma_{\sigma})n_{i,\sigma}$ (1)
where $\delta^{i}_{\sigma}=M_{z,\sigma}\cos{\left(2\pi\beta i\right)}$, with $M_{z,\sigma}\propto V_{s,\sigma}$ and $\gamma_{\sigma}$ denotes the incommensurate Zeeman potential strength and dissipation strength for spin $\sigma=\uparrow,\downarrow$, respectively. $n_{i,\sigma}=c^{\dagger}_{i,\sigma}c_{i,\sigma}$ is the particle number operator, $\beta$ is an irrational number and:
$H_{0}$ $=-t_{0}\sum_{i,\pm,\sigma}c_{i,\sigma}^{\dagger}c_{i\pm 1,\sigma}$ (2)
[…]
Here, $t_{0}$ ($t_{so}$) represents the spin-conserved (spin-flip) hopping strength between neighboring sites. The Hamiltonian renders a 1D AIII class topological insulator if replacing the incommensurate Zeeman potential with a constant uniform Zeeman potential Song et al. (2018); Liu et al. (2013); Wang et al. (2018); Zhang and Liu (2018).
Figure 1: Experimental realization of critical phase in optical Raman lattice with 173Yb atoms. (a) Experimental setup of the optical Raman lattice consists of a standard 1D optical Raman lattice $E_{1}$, $E_{3}$ with another spin-dependent incommensurate lattice $E_{21}$ and $E_{22}$. (b) Schematic energy diagram with relevant transitions. (c) The incommensurate lattice induces a spin-dependent offset $V_{2,\sigma}(x)$ to the standard optical Raman lattice potential. (d) For the implementation of a perfect spin-dependent incommensurate lattice, the laser detuning should be set approximately 1.9 GHz blue-detuned from the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition.
This Hamiltonian can be experimentally implemented through the configuration outlined in Fig. 1(a). To establish the conventional optical Raman lattice potential Liu et al. (2013); Wu et al. (2016); Song et al. (2018), a standing-wave beam $E_{1}$ with $y$ polarization and a free-running beam $E_{3}$ with z polarization are employed. The single-photon detuning of these two beams can be finely tuned to approximately 11 GHz blue-detuned from the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition, as depicted in Fig. 1(b). In this setup, the primary lattice, characterized by a wave vector of $k_{0}=2\pi/\lambda$, exhibits nearly spin-independent lattice potential $V_{1,\sigma}(x)=V_{p,\sigma}\cos^{2}{k_{0}x}$ with $V_{p,\uparrow}/V_{p,\downarrow}\sim 1.01$, where $V_{p,\uparrow}/V_{p,\downarrow}$ denotes the primary lattice depth for spin up or spin down atoms. In the subsequent discussion, we will neglect the spin-dependent effects of the primary lattice potential ($V_{p,\uparrow}=V_{p,\downarrow}=V_{p}$). Moreover, two additional beams propagating at an angle of 52 degrees relative to the x direction are utilized to induce a spin-dependent incommensurate lattice featuring a wave vector of $k_{1}=k_{0}\cos{\theta}$, giving rise to a secondary incommensurate Zeeman potential $V_{2,\sigma}=V_{s,\sigma}\cos^{2}{k_{1}x}$ with $V_{s,\sigma}$ representing the incommensurate lattice depth, as illustrated in Fig. 1(c). The spin-dependence of this incommensurate lattice can be controlled through the detuning $\Delta$ of this lattice beam with respect to the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition. When $\Delta\sim 1.9~{}\text{GHz}$, the incommensurate lattice becomes entirely spin-dependent with $M_{z,\uparrow}/M_{z,\downarrow}=-1$ (Fig. 1(d)). To achieve a complete spin-independent configuration, the detuning of the incommensurate lattice can be set larger than 13GHz. The Raman coupling in this setup can be denoted by $\mathcal{M}(x)=M_{R}\cos{k_{0}x}$. The introduction of dissipation into this Raman lattice setup can be realized through adding a nearly resonant loss beam with wavelength close to ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition Zhao et al. (2025).
For simplicity, all phases along the y direction in these potentials are neglected during the calculation and the Hamiltonian governing this experimental setup can be expressed as:
$H$ $=\left[\frac{\hbar^{2}k_{x}^{2}}{2m}+V_{1}(x)\right]\otimes\mathds{1}+\sum_{\sigma=\uparrow,\downarrow}V_{2,\sigma}(x)|\sigma\rangle\langle\sigma|$ (3)
$+\mathcal{M}(x)\sigma_{x}+\frac{\delta}{2}\sigma_{z}+\sum_{\sigma=\uparrow,\downarrow}\frac{i}{2}\gamma_{\sigma}|\sigma\rangle\langle\sigma|$
[…]
III III. Localization behavior in the Hermitian regime
Except for the localized phase, recent theoretical investigations have predicted the existence of a third fundamental phase, called critical phases, positioned between the localized and extended phases within such a spin-orbit coupled optical lattice featuring an incommensurate Zeeman potential Wang et al. (2020); Song et al. (2018); Zhao et al. (2025). Noteworthy features of these critical phases encompass critical spectral statistics Geisel et al. (1991); Machida and Fujita (1986); Bertrand and García-García (2016), multifractal properties of wave functions Halsey et al. (1986); Mirlin et al. (2006); Dubertrand et al. (2014), and dynamical evolutions Hisashi and Shuji (1988); Ketzmerick et al. (1997); Larcher et al. (2009). Moreover, recent theoretical investigations have also unveiled the possibility of a quantum phase harboring coexisting localized, extended, and critical regions within a one-dimensional quasi-periodic model Wang et al. (2022).
Therefore, we first consider an entirely spin-dependent incommensurate lattice ($M_{z,\uparrow}/M_{z,\downarrow}=-1$) in the Hermitian regime ($\gamma_{\sigma}=0$), and find that a critical phase manifests in this configuration. The phase of this system can be characterized through the mean fractal dimension $\bar{\eta}$, as introduced by Wang et al. (2020):
$\bar{\eta}=-\frac{\lim_{L\rightarrow\infty}\ln{\left[(2L)^{-1}\sum^{2L}_{m=1}\sum^{L}_{j=1}\left(u^{4}_{m,j}+v^{4}_{m,j}\right)\right]}}{\ln{2L}}$ (4)
Here, $\bar{\eta}$ is averaged over all eigenstates, which can be represented as:
$|{\psi_{m}}\rangle=\sum_{j=1}^{L}\left(u_{m,j}c^{\dagger}_{j,\uparrow}+v_{m,j}c^{\dagger}_{j,\downarrow}\right)|{\text{vac}}\rangle$ (5)
It is established that, in extended (localized) states, the value of $\bar{\eta}$ tends towards 1 (0), while within the critical phase, $10<\bar{\eta}<$ 1.
In Fig.2(a), the mean fractal dimension is depicted as a function of $M_{z}$ and $t_{so}$, revealing a distinct region where $10<\bar{\eta}<$ 1, indicating the emergence of the critical phase. The phase boundary can be delineated by $M_{z}=2|t_{0}\pm t_{so}|$, where the symbol $+$ signifies the boundary between localization and critical phases, and -- signifies the boundary between extended and critical phases Wang et al. (2020). Fig. 2(b) shows the fractal dimension for various eigenenergies and $M_{z}$ when $t_{so}=0.3$, illustrating that all individual states within the critical region exhibit critical behavior. However, if the incommensurate lattice is entirely spin-independent ($M_{z,\uparrow}/M_{z,\downarrow}=1)$, the critical phase region disappears, as illustrated in Fig. 2(c). Although a region where $\bar{\eta}$ deviates significantly from 1 and 0 can be identified, a further examination of the fractal dimension $\eta$ of the corresponding eigenstates as a function of $M_{z}$ suggests an intermediate regime characterized by the coexistence of extended and localized states rather than a critical phase (Fig.2(d)).
Figure 2: Phase diagram of the optical Raman lattice system with additional incommensurate lattice. (a) The various phases, namely extended, critical, and localization phases, are distinguished based on the mean fractal dimension $\bar{\eta}$ when $M_{z,\uparrow}/M_{z,\downarrow}=-1$. $M_{z}$ and $t_{so}$ are presented in units of $t_{0}$. (b) Fractal dimension $\eta$ and energy of individual states for distinct values of $M_{z}$ with $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=-1$. (c) Phase diagram of mean fractal dimension $\bar{\eta}$ as a function of $M_{z}$ and $t_{so}$ when $M_{z,\uparrow}/M_{z,\downarrow}=1$. Other parameters are identical to those in b. (d) Fractal dimension $\eta$ and energy of individual states for distinct values of $M_{z}$ with $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=1$. All results are simulated with $L=1597$ and $\beta=987/1597$ in periodic boundary condition.
In the pursuit of experimentally detecting localization behavior in an optical lattice system, one common method involves observing the expansion dynamics, as conducted in various studies Billy et al. (2008); Roati et al. (2008); Lüschen et al. (2018). Fig. 3(a) shows simulated time evolutions of wave packets positioned in the extended, critical, and localized phase regions. These simulations are based on an initial wave packet configuration defined as $|{\psi_{j}(t=0)}\rangle=(\sqrt{\pi}a)^{-1/2}e^{-(j-j_{0})^{2}/2a^{2}}|{\uparrow}\rangle$, where $a$ represents the half-width, and the packet is centered at site $j_{0}$. Notably, the outcomes reveal distinctive behaviors: in the extended phase, the atom cloud exhibits ballistic expansion, while in the localized phase, occupancy remains predominantly restricted to the initial site. In the critical phase, a quasi-localization phenomenon is observed, with the atom cloud spreading to neighboring unit cells around the initial site, without fully occupying the entire system.
To further characterize the expansion dynamics, the mean square displacement can be employed as an observable, defined as Wang et al. (2020):
$W(t)=\left[\sum_{j,\sigma}(j-j_{0})^{2}\langle n_{j,\sigma}(t)\rangle\right]^{1/2}$ (6)
In Fig.3(b), the evolution of $W$ over time for various $M_{z}$ values, with $a=5$, is presented. In the extended phase, $W$ rapidly converges to a large and stable value, while in the localized phase, $W$ remains consistently small. Conversely, in the critical phase, $W$ exhibits gradual and slow growth, distinguishing it from both the extended and localized regimes. Furthermore, the determination of phases can be facilitated by the time-averaged observable $\bar{W}=\frac{1}{N_{t}}\sum_{m=1}^{N_{t}}W(m\Delta T)$, where $N_{t}=100$ and $\Delta t=20$, representing the average width of the wave packet throughout the temporal evolution, as depicted in Fig. 3(c).
Figure 3: Simulation of detecting critical phase with expansion dynamics. (a) Time evolution of the wave packet for the system located in the extended phase ($M_{z}=0.5$), critical phase ($M_{z}=2.0$), and localized phase ($M_{z}=3.5$). (b) Expansion dynamics of the wave packet, characterized by $W$, as a function of $M_{z}$, with $t_{so}=0.3$ and an initial width of $a=5$. (c) Time-averaged $\bar{W}$ as a function of $M_{z}$ with different initial width $a$. All results are simulated with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary condition.
In most localization behavior studies, a crucial consideration lies in the necessity for the initial state to populate only a few lattice sites to ensure clarity in distinguishing between different phases (Fig.3(c)). While experimental setups in ultracold atom investigations enable the detection and preparation of Gaussian wavepackets as initial states through techniques like a matter wave magnifierMurthy et al. (2014); Asteria et al. (2021) or a quantum gas microscope Bakr et al. (2009); Sherson et al. (2010); Miranda et al. (2015); Haller et al. (2015); Edge et al. (2015); Omran et al. (2015); Cheuk et al. (2016); Parsons et al. (2016); Yamamoto et al. (2016); Mitra et al. (2018); Yamamoto et al. (2020); Kwon et al. (2022), this can introduce complexity and inconvenience to the experimental arrangement. An intriguing observation emerges from numerical simulations represented in Fig. 3(a), indicating that in the extended phase, the presence of spin-down atoms is negligible when all initial states are occupied by spin-up atoms. Contrastingly, in the critical phase, clear evolution of spin-down atoms is observable over time. This can provide another way to detect the emergence of a critical phase in the current system.
In Fig.4(a), the spin polarization for systems where all lattice sites are initially uniformly occupied by spin up atoms is depicted for various phases. Within the extended phase, spin polarization exhibits minor oscillations with the value close to 1, while in critical and localized states, distinct oscillations occur, with the critical phase displaying larger oscillation amplitudes than the localized phase. Comparing with the distinct boundary between the extended and critical phases, the demarcation between the critical and localized phases, while not as pronounced, remains identifiable. Utilizing time-averaged spin polarization can aid in characterizing the emergence of the critical phase (Fig.3(c)-(d)). Notably, the time-averaged spin polarization tends towards zero in the critical phase, contrasting with higher values outside this phase.
Figure 4: Evolution of spin polarization for different phases. (a) Time evolution of the spin polarization for the system located in the extended phase ($M_{z}=0.5$), critical phase ($M_{z}=2.0$), and localized phase ($M_{z}=3.5$) with the initial state comprising spin-up atoms in a uniform distribution. (b) Spin dynamics as a function of $M_{z}$, with $t_{so}=0.3$, and an initial distribution in the spin up state uniform distribution. (c) Time-averaged spin polarization as a function of $M_{z}$ averaged over $t=0-200\text{s}$, which show minimal disparity between initial states of uniform distribution or Gaussian distribution. (d) Time-averaged spin polarization as a function of $M_{z}$ and $t_{so}$. All simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary conditions.
IV IV. Coexistence of different states
Returning to the spin-independent incommensurate lattice case, although the critical phase is absent in such a system, under specific conditions, extended and localized states can coexist, separated by a mobility edge. For a more general situation, where the incommensurate Zeeman potential exhibits a blend of spin-dependence and independence, a phase emerges encompassing coexisting localized, extended, and critical regions Wang et al. (2022). Illustrated in Fig.5(a) is the fractal dimension corresponding to various eigenenergies and $M_{z}$, revealing regions where three or two distinct types of eigenstates coexist, separated by mobility edges. To validate the coexistence of difference phases within this domain, we consider the case of $M_{z}=1.5$ as an example (refer to Fig.5(b)). Notably, as the system size increases, $\eta$ converges towards 0 for all states in zones I and IV, indicative of localization states, while tending towards 1 for all states in zone III, signifying extended states. Conversely, within zones II and V, the fractal dimension markedly deviates from 0 and 1 and remains relatively constant with system size, suggesting a critical state in these regions.
The time evolution of the wave packet under such circumstances, as depicted in Fig.5(c), showcases characteristics signatures of these three distinct phases concurrently. On the whole, the time evolution of the wave packet exhibits quasi-localization, highlighting the presence of critical states. However, the higher concentration of atom clouds in the vicinity of central sites, particularly for spin-down atoms, indicates the existence of localized states. Furthermore, the inset of Fig.5(c) illustrates instances of ballistic expansion behavior of localized states upon closer inspection of the color bar. Through simulations of time evolution, the coexistence of extended, critical, and localized states under such conditions is further substantiated.
Figure 5: Phase with coexisting localized, extended, and critical zones. (a) Fractal dimension $\eta$ and energy of individual states for distinct values of $M_{z}$ with $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=-2$ with $t_{so}=0.3$. Simulations are conducted with $L=1597$ and $\beta=987/1597$ under periodic boundary condition. (b) Fractal dimension of different eigenstates as a function of eigenenregy for different system size $L=233$ and $L=1597$ with $M_{z}=1.5$ and $t_{so}=0.3$. For convenience, only the region with $E<0$ is presented , which is symmetric to that with $E>0$. (c) Time evolution of the wave packet for the system located in the phase with coexisting localized, extended, and critical zones ($M_{z}=1.5$, $t_{so}=0.3$ and $M_{z,\uparrow}/M_{z,\downarrow}=-2$). Simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary condition.
V V. Suppression of critical phase in non-Hermitian Raman lattice
Recently, various non-Hermitian extensions of systems exhibiting Anderson localization have been explored, such as the non-reciprocal Aubry-André model Jiang et al. (2019) and the PT-symmetric Aubry-André model Yuce (2014); Liang et al. (2014); Hang et al. (2015); Harter et al. (2016); Zeng et al. (2017); Longhi (2019). Numerical studies have delved into the impact of nonreciprocal tunneling, physical gain and loss on parity-time symmetry breaking, the Hofstadter butterfly spectrum Longhi (2014); Yuce (2014); Liang et al. (2014); Harter et al. (2016), topological phase transitions Longhi (2019), and the localization characteristics of eigenstates Hatano and Nelson (1996); Shnerb and Nelson (1998); Hang et al. (2015); Zeng et al. (2017); Jiang et al. (2019); Zhou (2023). Given these insights, it becomes pertinent to investigate the effects of non-Hermiticity on our system.
Expanding our setup into the non-Hermitian regime involves introducing dissipation to one of the spin states ($\gamma_{\sigma}\neq 0$) through a nearly resonant loss beam Zhou (2023). We focus on the scenario with $M_{z,\uparrow}/M_{z,\downarrow}=-1$, which exhibits a critical phase in the Hermitian context. The phase diagram for the system with dissipation ($\gamma_{\downarrow}=-2$) is illustrated in Fig. 6(a), revealing a transition of critical phases in the Hermitian regime to extended or localized phases in non-Hermitian regime. Fig. 6(b) presents the mean fractal dimension $\bar{\eta}$ for varying dissipation values $\gamma_{\downarrow}$ with $t_{so}={1.0}$. The plateau characterizing the critical phase in the Hermitian setup vanishes upon the introduction of dissipation. For system with larger $M_{z}$, the mean fractal dimension decreases, indicative of a transformation of eigenstates into localized states, whereas for lower $M_{z}$ values, the mean fractal dimension rises, signifying a transition towards extended states.
Furthermore, Fig. 6(c) shows the fraction of eigenstates exhibiting different fractal dimensions, depicting a critical-extended transition. Employing a methodology akin to that in Fig. 5, it is observed that in the Hermitian regime, the fractal dimension remains stable, indicative of critical states in these regions. However, upon introducing dissipation ($\gamma_{\downarrow}=2$), $\eta$ converges towards 1 for all states when the system size increases, confirming the transformation of nearly all states into extended states and the suppression of the critical phase within the system due to the induced dissipation (Fig. 6(d)). This suppression of the critical phase is attributed to non-Hermiticity, as evidenced by the persistence of this effect even under conditions where $\gamma_{\uparrow}/\gamma_{\downarrow}=-1$.
Figure 6: Suppression of critical phase in non-Hermitian Raman lattice. (a) The phases diagram after inducing dissipation ($\gamma_{\downarrow}=-2$) to the system, distinguished based on the mean fractal dimension $\bar{\eta}$ when $M_{z,\uparrow}/M_{z,\downarrow}=-1$. $M_{z}$ and $t_{so}$ are presented in units of $t_{0}$. The dashed lines show the boundary between different phases in Hermitian regime. (b) Mean fractal dimension $\bar{\eta}$ for different dissipation $\gamma_{\downarrow}$ as a function of $M_{z}$ with $t_{so}={1.0}$. (c) Fraction of eigenstates with $\eta>0.8$, $\eta<0.3$ and $0.80.3<\eta<$ 0.8 respectively as a function of $\gamma_{\downarrow}$ with $t_{so}=1.0$ and $M_{z}=1.0$. (d) Fractal dimension of different eigenstates as a function of eigenenergy for different system size $L=233$ and $L=1597$ with $M_{z}=1.5$ and $t_{so}=0.3$. The top one is without dissipation and the bottom one is with dissipation. For convenience, only the region with $E<0$ is presented , which is symmetric to that with $E>0$. Unless otherwise specified, all simulations are conducted with $L=1597$ and $\beta=987/1597$ under periodic boundary condition.
VI VI. Summary
This study introduces a versatile Raman lattice configuration aimed at investigating the localization behavior for alkaline-earth-like atoms, exemplified by 173Yb. The experimental setup explores three distinct quasi-periodic Raman lattice configurations: an entirely spin-dependent incommensurate lattice in a Hermitian regime, a partially spin-dependent incommensurate lattice under Hermitian conditions, and a fully spin-dependent incommensurate lattice in a non-Hermitian regime. In the case of the entirely spin-dependent incommensurate lattice in the Hermitian regime, the emergence of a critical phase is observed, which can be detected by expansion dynamics and spin polarization evolution. Conversely, when the incommensurate lattice transitions into a combination of spin-dependent and spin-independent characteristics, a new phase with coexistence of extended, critical, and localized phases separated by mobility edges shows up. On the other hand, we also find that introducing dissipation into the system leads to the suppression of the critical phase due to non-Hermiticity within specific parameter spaces. This work aims to stimulate further experimental investigations into the interplay between the localization behavior and non-Hermiticity Zhou (2023); Jiang et al. (2019) in ultracold atoms.
VII ACKNOWLEDGMENTS
GBJ acknowledges support from the RGC through 16306922, 16302123, 16305024, N-HKUST636-22, and RFS2122-6S04.

### v2

@@
 Abstract
-We introduce a flexible Raman lattice system for alakaline-earth like atoms to investigate localization behaviors in a quasi-periodic lattice with controllable non-Hermiticity.
-We demonstrate that critical phases and mobility edges can arise by adjusting spin-dependence of the incommensurate potentials in the Hermitian regime.
-With non-Hermiticity introduced by spin-selective atom loss, we find the localization behaviour in this system can be suppressed by dissipation.
+We propose a flexible Raman lattice system for alkaline-earth-like atoms to theoretically investigate localization behaviors in a quasi-periodic lattice with controllable non-Hermiticity.
+Our analysis demonstrates that critical phases and mobility edges can arise by adjusting spin-dependence of the incommensurate potentials in the Hermitian regime.
+With non-Hermiticity introduced by spin-selective atom loss, our calculations reveal that critical localization behaviour in this system can be suppressed by dissipation.
 Our work provides insights into interplay between quasi-periodicity and non-Hermitian physics, offering a new perspective on localization phenomena.
-I I. Introduction
-In the presence of a disorder potential, quantum systems can exhibit exponential localization of wave functions induced by wave interference, a phenomenon known as Anderson localization Anderson (1958).
+I Introduction
+In the presence of a disorder potential, quantum systems can exhibit exponential localization of wave functions induced by wave interference, a phenomenon known as Anderson localization [4].
 Ultracold atom systems, with their exceptional purity and high controllability, enable detailed exploration of localization phenomena.
-Recent experiments have investigated Anderson localization in optical lattices with either random disorder Billy et al.
-(2008) or incommensurate quasi-periodic structures Roati et al.
-(2008); Lüschen et al.
-(2018); Xiao et al.
-(2021).
+Recent experiments have investigated Anderson localization in optical lattices with either random disorder [11] or incommensurate quasi-periodic structures [59, 45, 74].
 For ultracold atoms, quasi-periodic optical lattices are created by superimposing two optical lattices with incommensurate wavelengths.
-In such setups, a notable transition from extended to localized states, commonly referred to as the Anderson transition, can manifest even within one-dimensional configurations Grempel et al.
-(1982); Aulbach et al.
-(2004), contrary to uncorrelated disordered systems where such transitions are expected to occur only in dimensions higher than two according to the scaling theory Thouless (1974); Abrahams et al.
-(1979).
-These systems also allow for the engineering of mobility edges-boundaries that separate coexisting localized and extended states at discrete energy levels Sarma et al.
-(1988); Evers and Mirlin (2008); Biddle and Das Sarma (2010); Lüschen et al.
-(2018); An et al.
-(2018, 2021).
-Recent theoretical studies suggest even more complex behavior, including intermediate critical phases and hybrid phases where localized, extended, and critical states coexist in specific one-dimensional quasi-periodic models Wang et al.
-(2020, 2022).
-Recent years have seen growing interest in non-Hermitian physics El-Ganainy et al.
-(2018); Ashida et al.
-(2021) across multiple research domains, including electrical circuits Helbig et al.
-(2020); Ezawa (2019), optics and photonics Lee et al.
-(2009); Xiao et al.
-(2020); Zhou et al.
-(2018); Weidemann et al.
-(2020), optomechanics Patil et al.
-(2022), NV centers Zhang et al.
-(2021); Yu et al.
-(2022), trapped ions Cao et al.
-(2023); Zhang et al.
-(2025), superconducting qubits Chen et al.
-(2021) and ultracold atom systems Li et al.
-(2019); Liang et al.
-(2022); Ren et al.
-(2022).
-Non-Hermitian systems display unique features such as exceptional points or rings Zhen et al.
-(2015); Ren et al.
-(2022); Liu et al.
-(2025), novel topological invariants, and the non-Hermitian skin effect Liang et al.
-(2022); Zhao et al.
-(2025).
-The Hatano-Nelson model, a fundamental non-Hermitian topological model, combines asymmetric hopping amplitudes with an on-site random potential Hatano and Nelson (1996).
-However, a comprehensive understanding of the interplay between disorder and non-Hermiticity, particularly the influences of dissipation on localization behavior, remains largely unexplored Zeng et al.
-(2017); Zhang et al.
-(2020); Harter et al.
-(2016).
+In such setups, a notable transition from extended to localized states, commonly referred to as the Anderson transition, can manifest even within one-dimensional configurations [21, 7], contrary to uncorrelated disordered systems where such transitions are expected to occur only in dimensions higher than two according to the scaling theory [66, 1].
+These systems also allow for the engineering of mobility edges-boundaries that separate coexisting localized and extended states at discrete energy levels [60, 18, 10, 45, 2, 3, 52].
+Recent theoretical studies suggest even more complex behavior, including intermediate critical phases and hybrid phases where localized, extended, and critical states coexist in specific one-dimensional quasi-periodic models [69, 70, 93, 94, 68].
+Recent years have seen growing interest in non-Hermitian physics [17, 5] across multiple research domains, including electrical circuits [28, 19], optics and photonics [35, 73, 89, 71], optomechanics [57], NV centers [84, 77], trapped ions [12, 85], superconducting qubits [13] and ultracold atom systems [36, 40, 58].
+Non-Hermitian systems display unique features such as exceptional points or rings [88, 58, 42], novel topological invariants, restoration of the dark state [92], and the non-Hermitian skin effect [40, 87, 50].
+The Hatano-Nelson model, a fundamental non-Hermitian topological model, combines asymmetric hopping amplitudes with an on-site random potential [26].
+However, a comprehensive understanding of the interplay between disorder and non-Hermiticity, particularly the influences of dissipation on localization behavior, remains largely unexplored [79, 80, 25].
 Furthermore, experimental methods for studying this relationship between disorder and non-Hermiticity remain limited.
-In this paper, we introduce a flexible Raman lattice scheme Liu et al.
-(2013) tailored for 173Yb atoms Song et al.
-(2018, 2019) or alkaline-earth metal atoms Liang et al.
-(2023) to investigate localization behaviors.
+In this paper, we propose a flexible Raman lattice scheme [41] tailored for 173Yb atoms [65, 64, 86] or alkaline-earth metal atoms [39] to theoretically investigate localization behaviors.
 The Raman lattice configuration induces quasi-periodicity using two optical lattices with incommensurate wavelengths, with the weak lattice potential akin to an incommensurate Zeeman potential.
-By adjusting the frequency of the incommensurate lattice close to the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition, the spin-dependent nature of the lattice potential may lead to the emergence of critical phases or mobility edges within the system.
-Finally, we generalize the Raman lattice with incommensurate potentials into the non-Hermitian regime with dissipation induced by atom loss Zhao et al.
-(2025).
-The analysis reveals that dissipation exerts a significant influence on the localization behaviors within the system, offering a platform to investigate the intricate interplay between disorder potential and non-Hermitian physics.
-The subsequent sections of this paper are structured as follows: Sec.II presents the model Hamiltonian of the flexible Raman lattice and outlines its experimental realization scheme.
-In Sec.III-V, we delve into the impact of the spin-dependency of the incommensurate lattice and dissipation within the system.
+Our model shows that adjusting the frequency of the incommensurate lattice close to the $^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition enables spin-dependent control, which can theoretically lead to the emergence of critical phases or mobility edges within the system.
+We further extend the Raman lattice model with incommensurate potentials into the non-Hermitian regime with dissipation induced by atom loss [87].
+Theoretical analysis reveals that dissipation exerts a significant influence on the localization behaviors within the system, providing a platform to model the intricate interplay between disorder potential and non-Hermitian physics.
+The subsequent sections of this paper are structured as follows: Sec.II presents the theoretical model Hamiltonian of the flexible Raman lattice and outlines its implementation scheme.
+In Sec.III-VI, we theoretically analyze the impact of the spin-dependency of the incommensurate lattice and dissipation within the system.
 Our focus will encompass three distinct scenarios: a completely spin-dependent incommensurate lattice, a partially spin-dependent incommensurate lattice, and an incommensurate lattice coupled with dissipation.
-Finally, Sec.VI offers a concise summary of the findings and conclusions derived from this investigation.
-II II. Model Hamiltonian
+Finally, Sec.VII offers a concise summary of the theoretical findings and conclusions derived from this investigation.
+II Model Hamiltonian
 The Hamiltonian considered in this paper can be written as:
-$H=H_{0}+\sum_{i,\sigma=\uparrow,\downarrow}(\delta_{\sigma}^{i}+i\gamma_{\sigma})n_{i,\sigma}$ (1)
-where $\delta^{i}_{\sigma}=M_{z,\sigma}\cos{\left(2\pi\beta i\right)}$, with $M_{z,\sigma}\propto V_{s,\sigma}$ and $\gamma_{\sigma}$ denotes the incommensurate Zeeman potential strength and dissipation strength for spin $\sigma=\uparrow,\downarrow$, respectively.
-$n_{i,\sigma}=c^{\dagger}_{i,\sigma}c_{i,\sigma}$ is the particle number operator, $\beta$ is an irrational number and:
-$H_{0}$ $=-t_{0}\sum_{i,\pm,\sigma}c_{i,\sigma}^{\dagger}c_{i\pm 1,\sigma}$ (2)
-[…]
+$H=H_{0}+\sum_{j,\sigma=\uparrow,\downarrow}(\delta_{\sigma}^{j}+i\gamma_{\sigma})n_{j,\sigma}$ (1)
+where $\delta^{j}_{\sigma}=M_{z,\sigma}\cos{\left(2\pi\beta j\right)}$, with $M_{z,\sigma}\propto V_{s,\sigma}$ and $\gamma_{\sigma}$ denotes the incommensurate Zeeman potential strength and dissipation strength for spin $\sigma=\uparrow,\downarrow$, respectively.
+$n_{j,\sigma}=c^{\dagger}_{j,\sigma}c_{j,\sigma}$ is the particle number operator, $\beta$ is an irrational number and:
+$H_{0}$ $=-t_{0}\sum_{j,\sigma}\left[\left(c_{j,\uparrow}^{\dagger}c_{j+1,\uparrow}-c_{j,\downarrow}^{\dagger}c_{j+1,\downarrow}\right)+\text{H.c.}\right]$ (2)
+$+t_{\text{so}}\sum_{j}\left[\left(c_{j,\uparrow}^{\dagger}c_{j+1,\downarrow}-c_{j,\uparrow}^{\dagger}c_{j-1,\downarrow}\right)+\text{H.c.}\right]$
 Here, $t_{0}$ ($t_{so}$) represents the spin-conserved (spin-flip) hopping strength between neighboring sites.
-The Hamiltonian renders a 1D AIII class topological insulator if replacing the incommensurate Zeeman potential with a constant uniform Zeeman potential Song et al.
-(2018); Liu et al.
-(2013); Wang et al.
-(2018); Zhang and Liu (2018).
-Figure 1: Experimental realization of critical phase in optical Raman lattice with 173Yb atoms.
+The Hamiltonian renders a 1D AIII class topological insulator if replacing the incommensurate Zeeman potential with a constant uniform Zeeman potential [65, 41, 67, 82].
+Figure 1: Experimental implementation scheme of critical phase in optical Raman lattice with 173Yb atoms.
 (a) Experimental setup of the optical Raman lattice consists of a standard 1D optical Raman lattice $E_{1}$, $E_{3}$ with another spin-dependent incommensurate lattice $E_{21}$ and $E_{22}$.
@@
 (c) The incommensurate lattice induces a spin-dependent offset $V_{2,\sigma}(x)$ to the standard optical Raman lattice potential.
-(d) For the implementation of a perfect spin-dependent incommensurate lattice, the laser detuning should be set approximately 1.9 GHz blue-detuned from the ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition.
+(d) For the implementation of a perfect spin-dependent incommensurate lattice, the laser detuning should be set to approximately 1.9 GHz blue-detuned from the $^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition.
 This Hamiltonian can be experimentally implemented through the configuration outlined in Fig.
 1(a).
-To establish the conventional optical Raman lattice potential Liu et al.
-(2013); Wu et al.
-(2016); Song et al.
-(2018), a standing-wave beam $E_{1}$ with $y$ polarization and a free-running beam $E_{3}$ with z polarization are employed.
+To establish the conventional optical Raman lattice potential [41, 72, 65, 83, 33], a standing-wave beam $E_{1}$ with $y$ polarization and a free-running beam $E_{3}$ with z polarization are employed.
 The single-photon detuning of these two beams can be finely tuned to approximately 11 GHz blue-detuned from the $^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition, as depicted in Fig.
@@
 In the subsequent discussion, we will neglect the spin-dependent effects of the primary lattice potential ($V_{p,\uparrow}=V_{p,\downarrow}=V_{p}$).
-Moreover, two additional beams propagating at an angle of 52 degrees relative to the x direction are utilized to induce a spin-dependent incommensurate lattice featuring a wave vector of $k_{1}=k_{0}\cos{\theta}$, giving rise to a secondary incommensurate Zeeman potential $V_{2,\sigma}=V_{s,\sigma}\cos^{2}{k_{1}x}$ with $V_{s,\sigma}$ representing the incommensurate lattice depth, as illustrated in Fig.
+Moreover, two additional beams propagating at an angle of $\pm$52 degrees relative to the x direction are utilized to induce a spin-dependent incommensurate lattice featuring a wave vector of $k_{1}=k_{0}\cos{\theta}$, giving rise to a secondary incommensurate Zeeman potential $V_{2,\sigma}=V_{s,\sigma}\cos^{2}{k_{1}x}$ with $V_{s,\sigma}$ representing the incommensurate lattice depth, as illustrated in Fig.
 1(c).
@@
 The Raman coupling in this setup can be denoted by $\mathcal{M}(x)=M_{R}\cos{k_{0}x}$.
-The introduction of dissipation into this Raman lattice setup can be realized through adding a nearly resonant loss beam with wavelength close to ${}^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition Zhao et al.
-(2025).
+The introduction of dissipation into this Raman lattice setup can be realized through adding a nearly resonant loss beam with wavelength close to $^{1}S_{0}(F=5/2)\rightarrow{}^{3}P_{1}(F=7/2)$ transition [87].
 For simplicity, all phases along the y direction in these potentials are neglected during the calculation and the Hamiltonian governing this experimental setup can be expressed as:
@@
 $+\mathcal{M}(x)\sigma_{x}+\frac{\delta}{2}\sigma_{z}+\sum_{\sigma=\uparrow,\downarrow}\frac{i}{2}\gamma_{\sigma}|\sigma\rangle\langle\sigma|$
-[…]
-III III. Localization behavior in the Hermitian regime
-Except for the localized phase, recent theoretical investigations have predicted the existence of a third fundamental phase, called critical phases, positioned between the localized and extended phases within such a spin-orbit coupled optical lattice featuring an incommensurate Zeeman potential Wang et al.
-(2020); Song et al.
-(2018); Zhao et al.
-(2025).
-Noteworthy features of these critical phases encompass critical spectral statistics Geisel et al.
-(1991); Machida and Fujita (1986); Bertrand and García-García (2016), multifractal properties of wave functions Halsey et al.
-(1986); Mirlin et al.
-(2006); Dubertrand et al.
-(2014), and dynamical evolutions Hisashi and Shuji (1988); Ketzmerick et al.
-(1997); Larcher et al.
-(2009).
-Moreover, recent theoretical investigations have also unveiled the possibility of a quantum phase harboring coexisting localized, extended, and critical regions within a one-dimensional quasi-periodic model Wang et al.
-(2022).
+Here, $\frac{\hbar^{2}k_{x}^{2}}{2m}$ is the kinetic energy term and $\delta$ denotes the two-photon detuning.
+The incommensurate lattice introduces a weak spin-dependent energy offset characterized by the irrational number $\beta=\cos\theta\approx 0.6157$, where $\beta$ is determined by the tilted angle.
+In our proposed configuration, which employs $m_{F}$ states within the same hyperfine manifold as $|{\uparrow,\downarrow}\rangle$, the spin-dependence of the incommensurate lattice can be independently tuned by manipulating the detuning of the lattice beam.
+This independent controllability is advantageous, as it becomes challenging to simultaneously tune $\beta$ and maintain a spin-dependent incommensurate lattice when $\beta$ is set by the laser wavelength ratio.
+As a result, our approach significantly broadens the range of accessible $\beta$ values.
+Notably, our method differs from existing schemes reported in [69, 70, 94].
+Those prior proposals were designed for alkali-metal atoms and rely on counter-propagating laser beams with wavelengths near the D1 and D2 transitions, rendering such schemes incompatible with alkaline-earth atoms due to the lack of these transitions [27].
+In those setups, the incommensurability parameter $\beta$ is determined by the ratio of the laser wavelengths, inherently restricting it to a narrow, specific range.
+Furthermore, spin dependence of the incommensurate lattice is controlled via the polarization of the laser light, which requires the spin states to be encoded in different hyperfine states.
+This contrasts with our configuration and previously demonstrated experimental realizations, where spin states are defined using different $m_{F}$ states within the same hyperfine manifold.
+Our scheme is particularly suited for fermionic alkaline-earth atoms such as 173Yb, a system already demonstrated in optical Raman lattices [65, 64, 87].
+The fermionic nature of 173Yb, together with the potential SU(N) symmetry of alkaline-earth atoms, can lead to phenomena distinct from bosonic systems, especially in the many-body regime [51, 68, 69].
+III Localization behavior in the Hermitian regime
+Except for the localized phase, recent theoretical investigations have predicted the existence of a third fundamental phase, called critical phases, positioned between the localized and extended phases within such a spin-orbit coupled optical lattice featuring an incommensurate Zeeman potential [69, 65, 87, 94].
+Noteworthy features of these critical phases encompass critical spectral statistics [20, 46, 9], multifractal properties of wave functions [23, 48, 15], and dynamical evolutions [29, 31, 34].
+Moreover, recent theoretical investigations have also unveiled the possibility of a quantum phase harboring coexisting localized, extended, and critical regions within a one-dimensional quasi-periodic model [70].
 Therefore, we first consider an entirely spin-dependent incommensurate lattice ($M_{z,\uparrow}/M_{z,\downarrow}=-1$) in the Hermitian regime ($\gamma_{\sigma}=0$), and find that a critical phase manifests in this configuration.
-The phase of this system can be characterized through the mean fractal dimension $\bar{\eta}$, as introduced by Wang et al.
-(2020):
+The phase of this system can be characterized through the mean fractal dimension $\bar{\eta}$, as introduced by [69]:
 $\bar{\eta}=-\frac{\lim_{L\rightarrow\infty}\ln{\left[(2L)^{-1}\sum^{2L}_{m=1}\sum^{L}_{j=1}\left(u^{4}_{m,j}+v^{4}_{m,j}\right)\right]}}{\ln{2L}}$ (4)
@@
 $|{\psi_{m}}\rangle=\sum_{j=1}^{L}\left(u_{m,j}c^{\dagger}_{j,\uparrow}+v_{m,j}c^{\dagger}_{j,\downarrow}\right)|{\text{vac}}\rangle$ (5)
-It is established that, in extended (localized) states, the value of $\bar{\eta}$ tends towards 1 (0), while within the critical phase, $10<\bar{\eta}<$ 1.
-In Fig.2(a), the mean fractal dimension is depicted as a function of $M_{z}$ and $t_{so}$, revealing a distinct region where $10<\bar{\eta}<$ 1, indicating the emergence of the critical phase.
-The phase boundary can be delineated by $M_{z}=2|t_{0}\pm t_{so}|$, where the symbol $+$ signifies the boundary between localization and critical phases, and -- signifies the boundary between extended and critical phases Wang et al.
-(2020).
+It is established that, in extended (localized) states, the value of $\bar{\eta}$ tends towards 1 (0), while within the critical phase, $0<\bar{\eta}<1$.
+In Fig.2(a), the mean fractal dimension is depicted as a function of $M_{z}$ and $t_{so}$, revealing a distinct region where $0<\bar{\eta}<1$, indicating the emergence of the critical phase.
+The phase boundary can be delineated by $M_{z}=2|t_{0}\pm t_{so}|$, where the symbol ++ signifies the boundary between localization and critical phases, and - signifies the boundary between extended and critical phases [69].
 Fig.
@@
 All results are simulated with $L=1597$ and $\beta=987/1597$ in periodic boundary condition.
-In the pursuit of experimentally detecting localization behavior in an optical lattice system, one common method involves observing the expansion dynamics, as conducted in various studies Billy et al.
-(2008); Roati et al.
-(2008); Lüschen et al.
-(2018).
+In the pursuit of experimentally detecting localization behavior in an optical lattice system, one common method involves observing the expansion dynamics, as conducted in various studies [11, 59, 45].
 Fig.
-3(a) shows simulated time evolutions of wave packets positioned in the extended, critical, and localized phase regions.
+3(a) shows simulated time evolutions (in units of $\hbar/t_{0}$) of wave packets positioned in the extended, critical, and localized phase regions.
 These simulations are based on an initial wave packet configuration defined as $|{\psi_{j}(t=0)}\rangle=(\sqrt{\pi}a)^{-1/2}e^{-(j-j_{0})^{2}/2a^{2}}|{\uparrow}\rangle$, where $a$ represents the half-width, and the packet is centered at site $j_{0}$.
@@
 In the critical phase, a quasi-localization phenomenon is observed, with the atom cloud spreading to neighboring unit cells around the initial site, without fully occupying the entire system.
-To further characterize the expansion dynamics, the mean square displacement can be employed as an observable, defined as Wang et al. (2020):
+To further characterize the expansion dynamics, the mean square displacement can be employed as an observable, defined as [69]:
 $W(t)=\left[\sum_{j,\sigma}(j-j_{0})^{2}\langle n_{j,\sigma}(t)\rangle\right]^{1/2}$ (6)
@@
 All results are simulated with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary condition.
+IV Spin dynamics and density imbalance
 In most localization behavior studies, a crucial consideration lies in the necessity for the initial state to populate only a few lattice sites to ensure clarity in distinguishing between different phases (Fig.3(c)).
-While experimental setups in ultracold atom investigations enable the detection and preparation of Gaussian wavepackets as initial states through techniques like a matter wave magnifierMurthy et al.
-(2014); Asteria et al.
-(2021) or a quantum gas microscope Bakr et al.
-(2009); Sherson et al.
-(2010); Miranda et al.
-(2015); Haller et al.
-(2015); Edge et al.
-(2015); Omran et al.
-(2015); Cheuk et al.
-(2016); Parsons et al.
-(2016); Yamamoto et al.
-(2016); Mitra et al.
-(2018); Yamamoto et al.
-(2020); Kwon et al.
-(2022), this can introduce complexity and inconvenience to the experimental arrangement.
-An intriguing observation emerges from numerical simulations represented in Fig.
-3(a), indicating that in the extended phase, the presence of spin-down atoms is negligible when all initial states are occupied by spin-up atoms.
-Contrastingly, in the critical phase, clear evolution of spin-down atoms is observable over time.
-This can provide another way to detect the emergence of a critical phase in the current system.
-In Fig.4(a), the spin polarization for systems where all lattice sites are initially uniformly occupied by spin up atoms is depicted for various phases.
-Within the extended phase, spin polarization exhibits minor oscillations with the value close to 1, while in critical and localized states, distinct oscillations occur, with the critical phase displaying larger oscillation amplitudes than the localized phase.
-Comparing with the distinct boundary between the extended and critical phases, the demarcation between the critical and localized phases, while not as pronounced, remains identifiable.
-Utilizing time-averaged spin polarization can aid in characterizing the emergence of the critical phase (Fig.3(c)-(d)).
-Notably, the time-averaged spin polarization tends towards zero in the critical phase, contrasting with higher values outside this phase.
+While experimental setups in ultracold atom investigations enable the detection and preparation of Gaussian wavepackets as initial states through techniques like a matter wave magnifier [53, 6] or a quantum gas microscope [8, 62, 47, 22, 16, 55, 14, 56, 75, 49, 76, 32], this can introduce complexity and inconvenience to the experimental arrangement.
+To address this constraint, we also numerically simulate the evolution of spin dynamics and density imbalance across varied initial conditions, thereby offering complementary pathways to distinguish phases within the current system.
+We begin by examining the initial state in which all lattice sites are occupied by spin-up atoms ($|{\psi}\rangle=|{\uparrow\uparrow\uparrow\cdots\uparrow}\rangle$).
+In the single-particle picture, this state corresponds to a momentum distribution localized at $k=0$.
+Figure 4(a), presents the time evolution of the spin polarization $I_{p}=(n_{\uparrow}-n_{\downarrow})/(n_{\uparrow}+n_{\downarrow})$ under this configuration.
+In the extended phase, spin polarization exhibits minor oscillations with the value close to 1, while pronounced oscillations are observed in both the critical and localized phases.
+Utilizing time-averaged spin polarization can aid in characterizing the transition between the critical phase and extend phase (Fig.4 (b)).
+Specifically, it approaches zero in the critical phase but remains close to one in the extended phase.
+To elucidate the origin of the distinct spin dynamics across the extended–critical transition, we simulate the reconstructed momentum distribution during the evolution (Fig.4(c)–(d)).
+In the extended phase, the momentum distribution remains localized near $k=0$ throughout the evolution.
+By contrast, in the critical phase, the distribution broadens to momentum far from $k=0$.
+In the absence of a quasi-periodic lattice, the system is described by the Bloch Hamiltonian [82, 65, 87, 41]:
+$\mathcal{H}(k)=(m_{z}-2t_{0}\cos{k})\sigma_{z}+2t_{\text{so}}\sin{k}~\sigma_{y}$ (7)
+At $k=0$, the off-diagonal term vanishes, leading to the suppression of spin dynamics at this momentum.
+Notably, in the present system, the transition between the critical and localized phases coincides with the topological phase boundary [69].
+Away from the topological regime, the system exhibits detuned spin oscillations.
+Although distinctions in spin dynamics do exist between the critical and localized phases, these differences are less pronounced than the clear demarcation observed between the extended and critical phases.
+Consequently, spin dynamics may not serve as a reliable indicator for identifying the critical‑to‑localized phase transition.
 Figure 4: Evolution of spin polarization for different phases.
-(a) Time evolution of the spin polarization for the system located in the extended phase ($M_{z}=0.5$), critical phase ($M_{z}=2.0$), and localized phase ($M_{z}=3.5$) with the initial state comprising spin-up atoms in a uniform distribution.
-(b) Spin dynamics as a function of $M_{z}$, with $t_{so}=0.3$, and an initial distribution in the spin up state uniform distribution.
-(c) Time-averaged spin polarization as a function of $M_{z}$ averaged over $t=0-200\text{s}$, which show minimal disparity between initial states of uniform distribution or Gaussian distribution.
-(d) Time-averaged spin polarization as a function of $M_{z}$ and $t_{so}$.
+(a) Spin dynamics as a function of $M_{z}$, with $t_{so}=0.3$, and an initial distribution in the spin up state with uniform distribution.
+(b) Time-averaged spin polarization as a function of $M_{z}$ and $t_{\text{so}}$.
+(c, d) Corresponding reconstructed momentum distribution during the time evolution at extended phase ($M_{z}=0.5$) and critical phase ($M_{z}=2.0$).
 All simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary conditions.
-IV IV. Coexistence of different states
+To further probe the critical‑to‑localized phase transition, we examine an alternative initial state in which spin-up atoms half-fill the lattice, described by $|{\psi}\rangle=|{0\uparrow 0\uparrow 0\uparrow\cdots}\rangle$.
+This state can be experimentally prepared and detected using an optical superlattice  [61, 45].
+Figure 5 (a) and (b) display the time evolution of the density imbalance between even and odd sites, defined as $I_{d,\sigma}=(N_{e,\sigma}-N_{o,\sigma})/(N_{e}+N_{o})$, where $\sigma=\uparrow,\downarrow$ and $N_{e}(N_{o})$ denotes the atom number on even (odd) sites, in both the critical and localized regimes.
+After long-time evolution, the density imbalance for spin-up atoms remains nonzero in the localized regime, whereas it decays to zero with oscillation in the critical regime.
+Interestingly, spin-down atoms tend to occupy lattice sites of the opposite parity, as indicated by the opposite sign of their density imbalance.
+This behavior may originate from the fact that the spin-flip term in our Hamiltonian includes only nearest-neighbor couplings without on-site spin flipping.
+We can further define the overall density imbalance as $I_{d}=I_{d,\uparrow}-I_{d,\downarrow}$ and show its time average $\bar{I}_{d}$ over the interval t=0200t=0-200 in Fig.
+5 (c).
+A clear boundary emerges between the localized and critical phases.
+Combined with the time-averaged spin polarization $\bar{I}_{p}$ introduced earlier, both phase transitions can be clearly identified (Fig.
+5), even in systems lacking the ability to prepare or detect Gaussian wavepackets localized to only a few lattice sites.
+Figure 5: Evolution of density imbalance for different phases.
+(a, b) Time evolution of the density imbalance for spin-up (a) and spin-down (b) atoms in the critical phase ($M_{z}=2.0,t_{\text{so}}=0.3$) and the localized phase ($M_{z}=5.0,t_{\text{so}}=0.3$).
+(c) Time-averaged density imbalance as a function of $M_{z}$ and $t_{\text{so}}$.
+The markers indicate the parameter sets corresponding to the dynamics shown in (a) and (b).
+(d) Time-averaged spin polarization $\bar{I}_{p}$ and density imbalance $\bar{I}_{d}$ as a function of $M_{z}$ averaged over the time interval t=0200t=0-200.
+The expected phase transition position is marked by a vertical dashed line.
+All simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary conditions.
+V Coexistence of different states
 Returning to the spin-independent incommensurate lattice case, although the critical phase is absent in such a system, under specific conditions, extended and localized states can coexist, separated by a mobility edge.
-For a more general situation, where the incommensurate Zeeman potential exhibits a blend of spin-dependence and independence, a phase emerges encompassing coexisting localized, extended, and critical regions Wang et al.
-(2022).
+For a more general situation, where the incommensurate Zeeman potential exhibits a blend of spin-dependence and independence, a phase emerges encompassing coexisting localized, extended, and critical regions [70].
 Illustrated in Fig.6(a) is the fractal dimension corresponding to various eigenenergies and $M_{z}$, revealing regions where three or two distinct types of eigenstates coexist, separated by mobility edges.
-To validate the coexistence of difference phases within this domain, we consider the case of $M_{z}=1.5$ as an example (refer to Fig.5(b)).
+To validate the coexistence of different phases within this domain, we consider the case of $M_{z}=1.5$ as an example (refer to Fig.6(b)).
 Notably, as the system size increases, $\eta$ converges towards 0 for all states in zones I and IV, indicative of localization states, while tending towards 1 for all states in zone III, signifying extended states.
 Conversely, within zones II and V, the fractal dimension markedly deviates from 0 and 1 and remains relatively constant with system size, suggesting a critical state in these regions.
-The time evolution of the wave packet under such circumstances, as depicted in Fig.5(c), showcases characteristics signatures of these three distinct phases concurrently.
+The time evolution of the wave packet under such circumstances, as depicted in Fig.6(c), showcases characteristic signatures of these three distinct phases concurrently.
 On the whole, the time evolution of the wave packet exhibits quasi-localization, highlighting the presence of critical states.
@@
 Simulations are conducted with $L=1597$ and $\beta=987/1597$ under periodic boundary condition.
-(b) Fractal dimension of different eigenstates as a function of eigenenregy for different system size $L=233$ and $L=1597$ with $M_{z}=1.5$ and $t_{so}=0.3$.
+(b) Fractal dimension of different eigenstates as a function of eigenenergy for different system size $L=233$ and $L=1597$ with $M_{z}=1.5$ and $t_{so}=0.3$.
 For convenience, only the region with $E<0$ is presented , which is symmetric to that with $E>0$.
@@
 Simulations are conducted with $L=400$ and $\beta=\cos{52^{\circ}}$ under open boundary condition.
-V V. Suppression of critical phase in non-Hermitian Raman lattice
-Recently, various non-Hermitian extensions of systems exhibiting Anderson localization have been explored, such as the non-reciprocal Aubry-André model Jiang et al.
-(2019) and the PT-symmetric Aubry-André model Yuce (2014); Liang et al.
-(2014); Hang et al.
-(2015); Harter et al.
-(2016); Zeng et al.
-(2017); Longhi (2019).
-Numerical studies have delved into the impact of nonreciprocal tunneling, physical gain and loss on parity-time symmetry breaking, the Hofstadter butterfly spectrum Longhi (2014); Yuce (2014); Liang et al.
-(2014); Harter et al.
-(2016), topological phase transitions Longhi (2019), and the localization characteristics of eigenstates Hatano and Nelson (1996); Shnerb and Nelson (1998); Hang et al.
-(2015); Zeng et al.
-(2017); Jiang et al.
-(2019); Zhou (2023).
+VI Localization behavior in non-Hermitian regime
+Recently, various non-Hermitian extensions of systems exhibiting Anderson localization have been explored, such as the non-reciprocal Aubry-André model [30] and the PT-symmetric Aubry-André model [78, 38, 24, 25, 79, 44].
+Numerical studies have delved into the impact of nonreciprocal tunneling, physical gain and loss on parity-time symmetry breaking, the Hofstadter butterfly spectrum [43, 78, 38, 25], topological phase transitions [44], and the localization characteristics of eigenstates [26, 63, 24, 79, 30, 91].
 Given these insights, it becomes pertinent to investigate the effects of non-Hermiticity on our system.
-Expanding our setup into the non-Hermitian regime involves introducing dissipation to one of the spin states ($\gamma_{\sigma}\neq 0$) through a nearly resonant loss beam Zhou (2023).
+Expanding our setup into the non-Hermitian regime involves introducing spin-dependent dissipation to one of the spin states ($\gamma_{\sigma}\neq 0$) through a nearly resonant loss beam [91].
+In such system, one generally focus on the post-selected space, which corresponds to the remaining atoms.
+Under this condition, both the eigenspectrum and the time evolution can be well descibed by the effective non-Hermtian Hamiltonian [37, 5].
+Without loss of the generality, in the therotical simulation, we consider a system with balanced gain and loss term ($\gamma_{\uparrow}=-\gamma_{\downarrow}=\gamma$).
 We focus on the scenario with $M_{z,\uparrow}/M_{z,\downarrow}=-1$, which exhibits a critical phase in the Hermitian context.
-The phase diagram for the system with dissipation ($\gamma_{\downarrow}=-2$) is illustrated in Fig.
-6(a), revealing a transition of critical phases in the Hermitian regime to extended or localized phases in non-Hermitian regime.
-Fig.
-6(b) presents the mean fractal dimension $\bar{\eta}$ for varying dissipation values $\gamma_{\downarrow}$ with $t_{so}={1.0}$.
-The plateau characterizing the critical phase in the Hermitian setup vanishes upon the introduction of dissipation.
-For system with larger $M_{z}$, the mean fractal dimension decreases, indicative of a transformation of eigenstates into localized states, whereas for lower $M_{z}$ values, the mean fractal dimension rises, signifying a transition towards extended states.
-Furthermore, Fig.
-6(c) shows the fraction of eigenstates exhibiting different fractal dimensions, depicting a critical-extended transition.
-Employing a methodology akin to that in Fig.
-5, it is observed that in the Hermitian regime, the fractal dimension remains stable, indicative of critical states in these regions.
-However, upon introducing dissipation ($\gamma_{\downarrow}=2$), $\eta$ converges towards 1 for all states when the system size increases, confirming the transformation of nearly all states into extended states and the suppression of the critical phase within the system due to the induced dissipation (Fig.
-6(d)).
-This suppression of the critical phase is attributed to non-Hermiticity, as evidenced by the persistence of this effect even under conditions where $\gamma_{\uparrow}/\gamma_{\downarrow}=-1$.
+The mean fractal dimension for the system with non-Hermitian term ($\gamma=2$) is illustrated in Fig.
+7(a).
+The region with $0<\bar{\eta}<1$ vanishes, potentially indicating that the critical phases in the Hermitian regime is suppressed.
+For system with larger $M_{z}$, the mean fractal dimension decreases to a near zero value, indicative of a transformation of eigenstates into localized states, whereas for lower $M_{z}$ values, the mean fractal dimension rises to near one value, signifying a transformation towards extended states.
+To avoid the finite size effect, we further present the fraction of extended, critical and localized phases for varying dissipation values in Fig.
+7(b) with $t_{\text{so}}=0.5$.
+In the non-Hermitian regime ($\gamma\neq 0$), the critical phase is significantly suppressed, and the critical phase changed into a mixed phase with coexisting localized and extended states.
+Furthermore, we can also find as the non-Hermitian term increases, the pure localized and extended phases are also enlarged (Fig.
+7 (c)).
+The mechanism for the emergence of critical states in the Hermitian regime of this spinful model also provides insight for their suppression under non-Hermitian perturbations.
+As revealed in Refs.[93, 94], critical states in such systems arise when the on-site potential exhibits generalized incommensurate zeros in its $\sigma_{z}$ component.
+These zeros can be mapped to incommensurately distributed zeros in the hopping coefficients of an effective one-dimensional spinless model.
+The presence of such zeros effectively fragments the system into decoupled segments, forcing the delocalized wave functions to reorganize and develop the characteristic self-similar structure of critical states.
+However, the introduction of non-Hermitian terms disrupts the condition for generalized incommensurate zeros in the on-site potential.
+Consequently, the formation of critical states is invariably suppressed.
 Figure 7: Suppression of critical phase in non-Hermitian Raman lattice.
-(a) The phases diagram after inducing dissipation ($\gamma_{\downarrow}=-2$) to the system, distinguished based on the mean fractal dimension $\bar{\eta}$ when $M_{z,\uparrow}/M_{z,\downarrow}=-1$.
+(a) Mean fractal dimension $\bar{\eta}$ of the system under a non-Hermtian term ($\gamma=2$) when $M_{z,\uparrow}/M_{z,\downarrow}=-1$.
 $M_{z}$ and $t_{so}$ are presented in units of $t_{0}$.
 The dashed lines show the boundary between different phases in Hermitian regime.
-(b) Mean fractal dimension $\bar{\eta}$ for different dissipation $\gamma_{\downarrow}$ as a function of $M_{z}$ with $t_{so}={1.0}$.
-(c) Fraction of eigenstates with $\eta>0.8$, $\eta<0.3$ and $0.80.3<\eta<$ 0.8 respectively as a function of $\gamma_{\downarrow}$ with $t_{so}=1.0$ and $M_{z}=1.0$.
-(d) Fractal dimension of different eigenstates as a function of eigenenergy for different system size $L=233$ and $L=1597$ with $M_{z}=1.5$ and $t_{so}=0.3$.
-The top one is without dissipation and the bottom one is with dissipation.
-For convenience, only the region with $E<0$ is presented , which is symmetric to that with $E>0$.
-Unless otherwise specified, all simulations are conducted with $L=1597$ and $\beta=987/1597$ under periodic boundary condition.
-VI VI. Summary
-This study introduces a versatile Raman lattice configuration aimed at investigating the localization behavior for alkaline-earth-like atoms, exemplified by 173Yb.
-The experimental setup explores three distinct quasi-periodic Raman lattice configurations: an entirely spin-dependent incommensurate lattice in a Hermitian regime, a partially spin-dependent incommensurate lattice under Hermitian conditions, and a fully spin-dependent incommensurate lattice in a non-Hermitian regime.
-In the case of the entirely spin-dependent incommensurate lattice in the Hermitian regime, the emergence of a critical phase is observed, which can be detected by expansion dynamics and spin polarization evolution.
-Conversely, when the incommensurate lattice transitions into a combination of spin-dependent and spin-independent characteristics, a new phase with coexistence of extended, critical, and localized phases separated by mobility edges shows up.
-On the other hand, we also find that introducing dissipation into the system leads to the suppression of the critical phase due to non-Hermiticity within specific parameter spaces.
-This work aims to stimulate further experimental investigations into the interplay between the localization behavior and non-Hermiticity Zhou (2023); Jiang et al.
-(2019) in ultracold atoms.
-VII ACKNOWLEDGMENTS
-GBJ acknowledges support from the RGC through 16306922, 16302123, 16305024, N-HKUST636-22, and RFS2122-6S04.
+(b) Fraction of eigenstates belonging to the extended, critical, and localized phases as a function of $M_{z}$ for different values of $\gamma$ at $t_{\text{so}}=0.5$.
+In the non-Hermitian regime ($\gamma\neq 0$), the critical phase is significantly suppressed.
+(c) As the system enters the non-Hermitian regime, the critical phase evolves into a mixed phase with coexisting localized and extended states.
+The purely localized and extended phases are also enlarged with increasing non-Hermitian strength.
+VII Conclusion and outlook
+This study proposes a versatile Raman lattice configuration for theoretically investigating localization behavior for alkaline-earth-like atoms, using 173Yb as an example.
+Unlike previous approaches that primarily employed alkali-metal atoms [69, 70, 94], 173Yb represents a practically demonstrated fermionic system in optical Raman lattices [65, 64, 87].
+The fermionic nature of 173Yb, along with the potential SU(N) symmetry of alkaline-earth atoms, can give rise to phenomena distinct from bosonic systems, particularly in the many-body regime [51, 68, 69].
+In this theoretical framework, we explore three distinct quasi-periodic Raman lattice configurations: an entirely spin-dependent incommensurate lattice in a Hermitian regime, a partially spin-dependent incommensurate lattice under Hermitian conditions, and a fully spin-dependent incommensurate lattice in a non-Hermitian regime.
+In the case of the entirely spin-dependent incommensurate lattice in the Hermitian regime, our model predicts the emergence of a critical phase, which can be theoretically detected through signatures in expansion dynamics and spin polarization evolution.
+Conversely, when the incommensurate lattice transitions into a combination of spin-dependent and spin-independent characteristics, theoretical calculations reveal a new phase with coexistence of extended, critical, and localized phases separated by mobility edges.
+Furthermore, we demonstrate that introducing non-Hermitian dissipation disrupts the generalized incommensurate zeros and further suppresses the critical phase.
+This work bridges a gap by providing a practical experimental scheme for implementing and studying localization with fermionic alkaline-earth atoms in optical Raman lattices, directly addressing experimental capabilities.
+The proposed scheme also facilitates exploration of the interplay between localization behavior and non-Hermiticity, thereby laying a foundation for future experimental studies with ultracold atoms [91, 30].
+Acknowledgements.
+We thank Xin-Chi Zhou and Xiong-Jun Liu for fruitful discussions. GBJ acknowledges support from the RGC through 16302123, 16305024 and RFS2122-6S04.
+Appendix A Finite size analysis
+Figure 8: Finite size effects on $N_{e},N_{c},N_{l}$ and $\eta$.
+(a)-(c) Fractal dimension $\eta$ of individual eigenstates computed at three system sizes $L$, corresponding to $M_{z}=1.0,2.0,3.0$ respectively.
+(d)-(f) Scaling of $\eta$ with system size $L$ across different zones.
+(g)-(i) Fractions of extended, critical, and localized states ($N_{e}/N,N_{c}/N,N_{l}/N$) as functions of system size $L$.
+All results are simulated with $\gamma=1.0$ and $t_{\text{so}}=0.5$ in periodic boundary conditions.
+In this section, we investigate the finite size effects on the fractal dimension and the fractions of extended, critical, and localized states ($N_{e}/N,N_{c}/N,N_{l}/N$) in the non-Hermitian regime.
+Figure 8(a)-(c) shows the fractal dimension $\eta$ for $t_{\text{so}}=0.5$ and $\gamma=1.0$ at different system sizes $L$.
+For lower $M_{z}$ (e.g., $M_{z}=1.0$), the fractal dimensions of all eigenstates increase with system size and approach 1, indicating extended behavior.
+Conversely, for higher $M_{z}$ (e.g., $M_{z}=3.0$), the fractal dimensions decrease with increasing $L$ and tend to zero, consistent with behaviors of localized states.
+In the intermediate regime ($M_{z}=2.0$), eigenstates with larger absolute eigenenergies ($|E|>1.15$) exhibit decreasing fractal dimensions as $L$ increases, while those with lower energies show the opposite trend, suggesting a phase with coexisting extended and localized states.
+Based on these observations, we classify eigenstates into distinct zones sharing similar localization properties.
+The average fractal dimension within each zone is presented in Figure 8(d)-(f).
+For $M_{z}=1.0$ and $3.0$, all eigenstates fall into a single zone, corresponding to a pure phase.
+At $M_{z}=2.0$, however, eigenstates separate into two zones with mean fractal dimensions evolving oppositely with system size—one increasing and the other decreasing.
+This allows us to assign consistent localization properties to all states within each zone and determine the fractions of extended, critical, and localized states.
+As shown in Figure 8(g)-(i), these fractions remain invariant with increasing system size.
+Appendix B Absence of non-Hermitian skin effect
+In this section, we validate the absence of the non-Hermitian skin effect in our system, confirming that the observed suppression of the critical phase does not originate from the skin effect.
+According to the established correspondence between winding numbers and skin modes in non-Hermitian systems [81, 87, 54, 90], a system exhibits the skin effect under open boundary conditions (OBC) if and only if the eigenvalue winding on the complex plane displays nontrivial closed-loop topology under periodic boundary conditions (PBC).
+The presence of the skin effect also leads to distinct eigenspectra under PBC and OBC.
+As shown in Figure 9, the eigenenergy spectra on the complex plane are presented for both boundary conditions.
+The PBC spectrum exhibits no loop structures, indicating the absence of the non-Hermitian skin effect in the present system.
+Figure 9: Eigenenergy spectrum on complex plane under PBC and OBC with $M_{z}=1.0,2.0,3.0$ respectively. All results are simulated with $\gamma=1.0$ and $t_{\text{so}}=0.5$.
+Appendix C Effect of the width of the initial wave packet on the expansion dynamics
+Figure 10: Effect of the width of initial wave packet.
+(a-d) Expansion dynamics of the wave packet, characterized by $W$ as a function of $M_{z}$, with $t_{so}=0.3$ and an initial width of $a=5,10,20,40$, respectively.
+(e) $W$ as a function of time $t$ for different widths with a fixed $M_{z}=0.5$.
+Other parameters are the same as the main Fig.
+3.
+Figure 11: Effect of the total evolution time.
+(a) Time evolution of the spin polarization for the system located in the extended phase ($M_{z}$ = 0.5), critical phase ($M_{z}$ = 2.0), and localized phase ($M_{z}$ = 3.5) with the initial state comprising spin-up atoms in a uniform distribution.(b) Time-averaged spin polarization $\bar{I}_{p}$ as a function of $M_{z}$ averaged over different time interval $t_{f}=10,30,100,200$.(c) Time-averaged density imbalance $\bar{I}_{d}$ as a function of $M_{z}$ averaged over different time interval $t_{f}=10,30,100,200$.
+Other parameters are identical to those in Fig.5(d).
+In this section, we investigate the effect of the initial wave packet width $a$ on the expansion dynamics.
+We find that while the initial wave packet width $a$ does not alter the qualitative distinction in dynamical behaviors among different phases, it does quantitatively influence the clarity of the contrast and the expansion speed, as shown in Figure 10.
+Appendix D Effect of the evolution time on time average $\bar{I}_{d}$ and $\bar{I}_{p}$
+In this section, we examine the influence of the total evolution time on the dynamics of the density imbalance and spin polarization.
+We observe that while a very short evolution time can quantitatively affect the time-averaged spin polarization and density imbalance, the qualitative trends across different phases remain unchanged, as demonstrated in Figure 11.
+Moreover, the time-averaged spin polarization is robust against the total evolution time once it spans several oscillation periods.
+Hence, despite experimental constraints that may limit the coherent evolution time, spin dynamics can still serve as a reliable detection scheme for distinguishing between phases—provided the total evolution time is sufficiently long to encompass several periods of spin oscillation.